OpenStudy (anonymous):

simplify the expression: 2sin^2x-1/sinx-cosx Can someone please explain all the steps...i'm really confused...

OpenStudy (anonymous):

\[2\sin ^{2}x-1\div sinx-cosx\] here it is in equation form

OpenStudy (anonymous):

Use the fact that: \[\cos^2x-\sin^2x=1-2\sin^2x\]

OpenStudy (zepp):

http://puu.sh/qjVR

OpenStudy (zepp):

http://puu.sh/qjXJ Hope you can understand better with this ;D

OpenStudy (anonymous):

@zepp ty so much ^^ i understand now

OpenStudy (anonymous):

\[\cos(2*x)=(\cos(x))^{2}-(\sin(x))^{2}\rightarrow 1-2*(\sin(x))^{2}=\cos(2*x)\] your equation will be: \[-\cos(2*x)/(\sin(x)-\cos(x))\] multiplying all by \[\sin(x)+\cos(x)\] we have: \[-[\cos(2*x)/(\sin(x)^{2}-\cos(x)^{2})]*(\sin(x)+\cos(x))=\sin(x)+\cos(x)\]

OpenStudy (anonymous):

@RaphaelFilgueiras ty ^^

OpenStudy (zepp):

Haha, you're welcome ;D Good luck on simplifying other fractions! ^_^